The road was Rocky, so after running for awhile at 10 km/h Clytemnestra slowed to a five kilometer per hour walk if the total trip was 65 kilometers and she made it in 9 hours how far did she walk and how far did she run?
Answer by mananth(16946)


running speed = 10km/h
running distance = x km
running time = x/10
walking speed = 5 km/h
walking distance = 65-x
walking time = 65-x/5
walking time + running time = 9
(65-x)/5 + x/10 =9
multiply equation ny 110
2(65-x)+x =90
130-2x+x=90
x=40km running distance